OpenStudy (anonymous):

Line M is represented by the following equation: x – y = 8 Which equation completes the system that is satisfied by the solution (18, 10)? (4 points) x – y = –28 2x – 2y = 36 x + y = 18 2x – y = 26

OpenStudy (anonymous):

@Saber2th17

OpenStudy (bibby):

plug (18,10) into each equation until you get an answer of the form x = x for example, 2x-2y=36 2(18)-2(10)=36 36-20=36 16=36 so not the answer

OpenStudy (anonymous):

okay...